A box of volume 1 cm3 contains 5.2×1021 electrons. Calculate their Fermi

momentum and Fermi energy. Take me=9.1*10-28g ,mn=1.67×10-24g 1.67×10-27 ergs.


1
Expert's answer
2021-03-14T19:15:06-0400

The Fermy energy can be calculated as follows:


"E_F=\dfrac{h^2}{2m_e}(\dfrac{3N}{8\pi V})^{2/3},""E_F=\dfrac{(6.62\cdot10^{-34}\ J\cdot s)^2}{2\cdot9.1\cdot10^{-31}\ kg}\cdot(\dfrac{3\cdot5.2\cdot10^{21}\ el}{8\pi\cdot1\cdot10^{-6}\ m^3})^{2/3}=1.75\cdot10^{-19}\ J."

By the definition of the Fermi momentum, we have:


"p_F=\sqrt{2m_eE_F}=\sqrt{2\cdot9.1\cdot10^{-31}\ kg\cdot1.75\cdot10^{-19}\ J}=5.64\cdot10^{-25}\ \dfrac{kgm}{s}."